If a roof of a home is 7m from the peak to eave and the attic space is 3cm high at the peak, how long does each of the pieces of insulation need to be? Round to the nearest tenth.
Any help on this word problem would be great, thank you so much

Answer by checkley71(8403)

THE TRIANGLE FORMED IS 7M FOR THE HYPOTENUSE, .03M (3CM)FOR ONE SIDE AND X FOR THE UNKNOWN SIDE.
7^2=.03^2+X^2
49=.0009+X^2
X^2=49-.0009
X^2=48.9991
X=SQRT48.9991
X=6.9994 METERS LONG
-----------------------------------------------
IF THESE FIGURES ARE CORRECT 7M & 3CM. YOU HAVE NO ROOM TO INSTALL THE INSULATION.
####################################################################
HAPPY NEW YEARS